the following data given a paired yeilds of two varities that the mean yeilds are equal
Varity A : 45,32,58,57, 60,38,47,51,42,38
Varity B : 47,34,60,59,63,44,49,53,46,41
1
Expert's answer
2021-08-11T19:28:24-0400

"H_0:\mu_d=0"

"H_1:\mu_d \ne0"


"t=\frac{\mu_d}{\frac{sd}{\sqrt{n}}}"

"t=\frac{-2.8}{\frac{1.316561}{\sqrt{10}}}"

"=-6.7254"

"CV=T_{0.05,9}=\pm2.2622" (Two tailed)

Reject the null hypothesis since, the absolute value of the test statistic is greater than the absolute value of the critical value. There is a significant difference between mean yields.